Write an exponential function y=ab^x whose graph passes through the given points: (2,16), (3,64)
Answer by scott8148(6628)

substituting __ 16=ab^2

substituting __ 64=ab^3

dividing __ 64/16=(ab^3)/(ab^2) __ 4=b

substituting __ 16=a(4^2) __ 1=a